OpenStudy (anonymous):

whats the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (aroub):

-3y=-5x-6 y=5/3x+2 slope=5/3 y-intercept=2

OpenStudy (anonymous):

Solve for y. The slope if the coefficient on the x variable. The intercept is the solution you get for x=0

OpenStudy (anonymous):

Or you can just look at the 'b' term

OpenStudy (amistre64):

given the form: ax + by = c, slope is ultimately -a/b

OpenStudy (amistre64):

also given the form: ax + by = c , the intercepts can be read as \[\frac{x}{c/a}+\frac{y}{c/b}=1\] such that the x int is: c/a and the y int is: c/b

OpenStudy (anonymous):

Always gotta be makin up formulae.

OpenStudy (amistre64):

its a hobby :)